OpenStudy (anonymous):

Find the polynomial function with roots 1, 7, and –3 of multiplicity 2

OpenStudy (anonymous):

(x - 1)(x - 7)(x + 3)(x + 3) = 0 Can you multiply this out?

OpenStudy (anonymous):

you mean foil it?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

how would i foil 4 terms?

OpenStudy (anonymous):

Do by pair: (x + 3)(x + 3)

OpenStudy (anonymous):

ohh okay. x^2+3x+3x+9

OpenStudy (anonymous):

and the other x^2+7x+7x+7

OpenStudy (anonymous):

i messed up sorry .

OpenStudy (anonymous):

x^2-7x-x+7

OpenStudy (anonymous):

x^2+3x+3x+9 = x² + 6x + 9 Then multiply ( x + 7 ) ( x² + 6x + 9 )

OpenStudy (anonymous):

–3 of multiplicity 2 <=> ( x +3)²

OpenStudy (anonymous):

Alright. After combining like terms, you get \((x^{2} + 6x + 9)(x^{2} - 8x + 7)\). Now, in order to multiply trinomials, you use the same concept of FOILing, but maybe this'll help. What I have ^ turns to: \[x^{2}(x^{2}) + x^{2}(-8x) + x^{2}(7) + 6x(x^{2}) + 6x(-8x) + 6x(7) + 9(x^{2}) + 9(-8x) + 9(7)\]Can you solve this and combine like terms?

OpenStudy (anonymous):

\[\small x^{2}(x^{2}) + x^{2}(-8x) + x^{2}(7) + 6x(x^{2}) + 6x(-8x) + 6x(7) + 9(x^{2}) + 9(-8x) + 9(7)\]

OpenStudy (anonymous):

x^4-8x^2+6x^2-48x^2+42x+9x^2-72x+63

OpenStudy (anonymous):

How'd you get that? Remember the exponent rules. \((x^{a})(x^{b}) = x^{a + b}\)

OpenStudy (anonymous):

@Calcmathlete Let me show the easier way

OpenStudy (anonymous):

Alright. You're gonna do it a binomial at a time like you were about to before?

OpenStudy (anonymous):

yeah, I think the asker stuck with 3 terms

OpenStudy (anonymous):

Yeah. it's just expanding the concept of FOILing, so I leave the rest to you :)

OpenStudy (anonymous):

Can you multiply ( x + 7 ) ( x² + 6x + 9 ) = ...

OpenStudy (anonymous):

( x³ + 6x² + 9x ) + ( 7x² + 42x + 63) Can you add them together?

OpenStudy (anonymous):

Then take the result multiply for ( x -1)

OpenStudy (anonymous):

@Jarelii Can you try?

OpenStudy (anonymous):

sorry but this just really hard i give up i dont get it